OpenStudy (anonymous):

Solve: 5(x + 8) = 7(x − 2) A.17 B.−17 C.27 D.−27

hartnn (hartnn):

distribute using a(b+c)=ab+ac

hartnn (hartnn):

u get 5x+40 = 7x-14 can u solve further ??

OpenStudy (nubeer):

it's simple multiplication .. multiply 5 with x and 8 .... and multiply 7 with x and -2

OpenStudy (anonymous):

as hartnn mentioned ,a(b+c)=ab+ac implies, 5(x+8)=5*x + 5*8 = 5x+40 and 7(x-2)=7*x-7*2=7x-14 = 5(x+8)=7(x-2) = 5x+40=7x-14 now subtract both sides by 5x, = 5x+40-5x=7x-14-5x = 40 = 2x-14 now add 14 both sides, = 40+14 = 2x-14+14 = 54 = 2x now divide both sides by 2 , = 54/2 = 2x/2 = 37=x

OpenStudy (anonymous):

this is so confusing!

OpenStudy (nubeer):

hmm well 3 of us explained in our own way.. can u tell on which specific part u r stuck at?

OpenStudy (anonymous):

well I thinkk shubham.bagrecha explained it the best out of all. I am trying to look at it from his point, but its still confusing.